Scott travelled at a constant speed of 4.8 m/s. He had a mass of 83 kg. What was his kinetic energy?


1
Expert's answer
2021-11-04T10:16:23-0400
"KE=\dfrac{1}{2}mv^2,""KE=\dfrac{1}{2}\times83\ kg\times(4.8\ \dfrac{m}{s})^2=956\ J."